OpenStudy (zzr0ck3r):

Let \(g\in C([0,1])\). Consider the linear operator \(L_g:C([0,1])\rightarrow C([0,1])\) defined by \(L_g(f) = fg\). Show that \(L_g\) is continuous and find \(|||L_g|||\).

OpenStudy (zzr0ck3r):

I did the continuous part, but now I need to find \(|||L_g|||\). @eliassaab

OpenStudy (anonymous):

ah, linear operators. i remember learning linear/abstract algebra as if it were yesterday.

OpenStudy (kainui):

Tell us more @bloopman

OpenStudy (zzr0ck3r):

it was for me

OpenStudy (zzr0ck3r):

but this is for real analysis:)

OpenStudy (anonymous):

unfortunately i'm busy with other work of my own right now, but i'll try to get to a good stopping point to try to help

OpenStudy (kainui):

I feel like this hinges on the definition of continuity.

OpenStudy (anonymous):

oh. i've never looked into R.A.; just proofs it seems. either way, this is part of abstract algebra, which is how i learned linear operators anyway i'll get back to work

OpenStudy (zzr0ck3r):

\(|||L||| = \sup\{ \ ||L(f)|| \ : ||f||\le 1\}\)

OpenStudy (zzr0ck3r):

well because its continuous we know \(|||L|||<\infty\)

OpenStudy (anonymous):

Notice that, let f be such \( || f|| \le 1\) \[ ||| L_g|||\le || g f||= \sup_{0\le t \le 1}| g(t) f(t)| \le\sup_{0\le t \le 1}| g(t)|=||g||\\ ||| L_g||| \ge ||L_g(1)||=||g||\\ ||| L_g||| =||g|| \]

OpenStudy (zzr0ck3r):

ty

OpenStudy (anonymous):

YW